OpenStudy (3psilon):

The toll for trucks to cross a certain bridge is a total of $2 for the first two axles plus $2 for each additional axle. What function would give the toll , T(n) , in dollars, for a truck with n axles.

OpenStudy (3psilon):

I got T(n) = 2n+2 but that isn't right! Why?

OpenStudy (anonymous):

T(n)=2+(n-2)*2=2+2n-4=2n-2

OpenStudy (3psilon):

Where did you get n-2 @sauravshakya ?

OpenStudy (anonymous):

There are n no. of axles.... for the first two it cost 2...... Now, there are (n-2) remaining axles which cost 2 per axle

OpenStudy (anonymous):

got it?

OpenStudy (3psilon):

Ohhhhhh That makes a lot of sense! Thanks!

OpenStudy (anonymous):

WELCOME